# ReNiL: Relative Neural Inertial Locator with Any-Scale Bayesian Inference

### v2

@@ Abstract
Pedestrian inertial localization plays a crucial role in mobile and Internet of Things (IoT) services, offering ubiquitous positioning without requiring external infrastructure. However, existing deep-learning-based pedestrian inertial localization methods rely on fixed sliding-window integration, lack flexibility in adapting to varying motion scales and cadences, and provide inconsistent uncertainty estimation. This limits their utility in real-world mobile computing scenarios. In this paper, we propose Relative Neural Inertial Locator (ReNiL), a novel Bayesian deep-learning framework for accurate, efficient, and uncertainty-aware pedestrian localization. ReNiL introduces the concept of Inertial Positioning Demand Points (IPDPs) and focuses estimation on contextually meaningful points rather than continuous dense tracking. It supports inference on inertial measurement unit (IMU) sequences at any scale, enabling flexible positioning cadences aligned with application needs. ReNiL integrates a motion-aware orientation filter with an Any-Scale Laplace Estimator (ASLE), a dual-task deep network that combines patch-based self-supervision and Bayesian regression. By parameterizing displacements with a Laplace distribution, ReNiL yields homogeneous Euclidean uncertainty estimates, thereby facilitating seamless fusion with other positioning sources. A Bayesian inference chain connects successive IPDPs to form an efficient and unified localization process. Extensive experiments on RoNIN-ds and our new WUDataset, which feature complex indoor/outdoor motion from 28 participants, have demonstrated that ReNiL achieves state-of-the-art displacement accuracy and uncertainty consistency, outperforms TLIO, CTIN, iMoT, and RoNIN variants, and reduces computational overhead. Application studies have confirmed ReNiL’s robustness and usability for mobile and IoT localization tasks, establishing it as a scalable, uncertainty-aware foundation for next-generation mobile positioning systems.
Index Terms:
@@
Pedestrian Inertial Localization, Inertial Positioning Demand Points, Bayesian Process, Any Scale, Deep Learning
I Introduction
The ubiquitous use of smart mobile devices has resulted in consistently growing demand for location-based services. Pedestrian localization has become a key area of research. Map-based localization methods (e.g., WiFi fingerprinting [1], BLE beacons [2], UWB [3], and RFID [4, 5]) provide absolute position references, while visual inertial odometry (VIO) [6] and laser inertial odometry (LIO) [7] provide high-precision relative estimation. However, they require hardware installation, map construction, or rely on additional devices, and raise privacy concerns. In contrast, an inertial navigation system (INS) performs relative positioning by integrating data from IMU. This enables it to be used in any scenario that complies with Newtonian mechanics, without requiring prior knowledge of the environment and avoiding privacy issues. The application of IMU in pedestrian localization systems attracts widespread attention, including trajectory reckoning, signal fusion, and graph optimization [8, 9].
Recent progress has been driven by the improvement of consumer-grade MEMS sensors and the adoption of deep learning techniques. This wave of research has begun with kinematic modeling based on Newtonian mechanics, has progressed to data-driven displacement estimation, and has evolved into methods that support fusion with other positioning sources. Newtonian kinematic modeling is theoretically consistent with applied physics, but it performs poorly on consumer-grade IMUs [10, 11]. Data-driven methods show that well-trained neural networks, including convolutional [12, 13, 14], recurrent [15, 16], and attention-based [17, 18] architectures, can accurately estimate inertial relative displacement from IMU sequences, outperforming traditional approaches on consumer devices. Some studies use probabilistic modeling of regression targets to improve performance and provide rough uncertainty estimates [13, 16, 15, 17], while others work on the fusion of specified external signals and IMU [19, 20, 21, 22].
Considering the practical demands of pedestrian location services and the inherent limitations of existing methods, the fundamental goal of a pedestrian localization system is to provide accurate position estimations and error analysis when positioning is required. However, current inertial navigation models predominantly focus on precise trajectory tracking over entire sequences, operating at fixed time intervals to output continuous position updates. This fixed update cadence overlooks the core requirement of users for context-relevant positioning and makes it difficult to estimate consistent uncertainty. We summarize the following limitations among existing inertial positioning solutions.
  • Demand-driven estimation at inertial positioning demand points (IPDPs). In practical pedestrian services, such as indoor positioning, geofencing alerts, context-aware content, position updates are required only at meaningful moments. Present inertial solutions stream dense fied-rate trajectories, incurring superfluous computation, energy cost, and error accumulation when the sampling cadence diverges from the application’s needs.
  • Application-specific probabilistic uncertainty inference. Every IPDP estimate must be accompanied by a calibrated error model that: (i) remains numerically stable across variable-length inertial segments; (ii) is expressed in Euclidean space to permit principled fusion with heterogeneous sources (e.g., GNSS covariances, map constraints); and (iii) supports chained error estimation and propagation. Existing deep inertial odometry either omits uncertainty or reports scale-dependent scores, hindering multi-source coupling and downstream error analysis.
These limitations necessitate a novel neural inertial localization framework that prioritizes the accurate estimation of relative displacement and uncertainty at IPDPs. This framework should accommodate IMU sequences of varying scales, flexibly adapt to different localization cadences, minimize computational cost, and provide uncertainty estimates consistent with Euclidean homogeneity, thereby ensuring high availability with or without sensor fusion applications.
@@ I Introduction
To this end, we propose ReNiL, a deep learning-based Bayesian inference framework for inertial localization. ReNiL incorporates a motion-aware orientation filter specifically designed for pedestrian dynamics, enhancing the alignment between IMU measurements and the navigation coordinate system. To support IMU sequences of any scale and enable uncertainty estimation, it employs a dual-task deep network, Any-Scale Laplace Estimator (ASLE), which combines patch-wise self-supervision with Bayesian regression. Furthermore, ReNiL establishes a principled Bayesian inference pipeline that forms a continuous chain of displacement and uncertainty estimates across Inertial Positioning Demand Points (IPDPs).
@@
When the user starts using the inertial service, the system collects IMU data to adaptively estimate the device orientation and gradually align the navigation coordinate system with the ENU coordinate system. ReNiL first aligns the entire IMU sequence with the navigation coordinate system based on the device orientation at each timestamp and then patches the data according to a given patch length. The patched data are processed in parallel by an embedding module and a feature extractor to extract motion features. To improve temporal correlation, the features are reconstructed into robust contextual motion features by a contextual builder. A scale-pooling mechanism is designed to build unified global features and contextual features, allowing it to regress to different scales. The self-supervision patch method is used to further improve the noise resistance of the model and the contextual learning capabilities. To handle potential aleatoric uncertainty in the data and improve the numerical stability of regression, displacement and its uncertainty estimates are parameterized using the Laplace distribution [23]. Additionally, ReNiL provides a general process based on Bayesian inference for pure IPDP tasks or tasks loosely coupled with external positioning sources.
@@
Extensive experiments on the open-source RoNIN-ds and our own WUDataset have demonstrated that ReNiL achieves state-of-the-art displacement accuracy and uncertainty consistency, surpassing all advanced models. We have verified the validity of parameterizing the displacements with the Laplace distribution and validated its Euclidean homogeneity. We have analyzed the computational overhead of ReNiL and several other models, and verified that ReNiL can greatly reduce the overhead of floating-point operations. We have also validated our proposed motion-aware orientation filter on WUDataset, and its accuracy in pedestrian motion far exceeds that of general methods. Finally, we have conducted a series of practical application experiments, demonstrating the high availability of ReNiL in real-world scenarios by combining the Bayesian process.
@@
  • IPDP task formulation. To the best of our knowledge, we firstly formulate the IPDP problem, reframing pedestrian localization as demand driven displacement and uncertainty estimation rather than fixed-rate odometry.
@@
  • Comprehensive validation. Experiments on RoNIN-ds and our WUDataset (will be released soon) show that ReNiL achieves state-of-the-art displacement accuracy, Euclidean homogeneous uncertainty, and markedly reduces FLOPs.
The paper is organized as follows. Section II reviews the relevant literature. Section III introduces the formulation and overall workflow of ReNiL. Section IV presents the detailed design of ReNiL. Section V provides experimental results. Section VI concludes the paper.
II Related Work
II-A Pedestrian Dead Reckoning
Pedestrian Dead-Reckoning (PDR) estimates a pedestrian’s position by integrating step detection, stride length estimation, and heading direction from inertial sensors [24]. Traditional PDR methods are highly dependent on heuristic models and sensor calibration, which may suffer from cumulative drift and noise. Recent work has focused on improving robustness by using map information [25] or combining additional signals such as magnetometers and barometers [26]. However, conventional PDR still faces challenges in complex indoor environments with magnetic disturbances and sensor biases [27].
@@ II-A Pedestrian Dead Reckoning
II-B Deep Learning for Inertial Navigation
Deep learning has been increasingly applied to inertial navigation to address the limitations of traditional algorithms. Data-driven methods learn representations from raw or aligned IMU data to directly estimate displacement. RNNs, especially LSTM networks, have shown effectiveness in modeling temporal dependencies in IMU sequences [15] [16]. CNN-based architectures have been explored to effectively suppress short-term drift in sensor data [12] [13] [14]. Attention-based architectures have also been explored to capture long-range correlations in sensor data [17] [18]. These methods reduce the need for manual calibration and demonstrate improved accuracy.
II-C Deep Learning for Time Series
Recent advancements in deep learning techniques for time series forecasting have been exemplified by models such as PatchTST, TimesNet, and TimeMixer. PatchTST [28], a Transformer-based architecture that represents time series data as patches, offering an efficient design for multivariate time series forecasting and self-supervised representation learning. TimesNet [29], a one-dimensional time series in a two-dimensional tensor that effectively captures intracycle and intercycle variations. This transformation allows the model to more accurately identify complex temporal patterns. TimeMixer [30], is capable of decomposing multiscale fusion, simultaneously capturing both short-term and long-term temporal dependencies.
II-D Deep Learning for Uncertainty
In the context of deep learning, uncertainty modeling has emerged as a critical technique to improve model reliability and robustness, particularly in domains such as computer vision and trajectory prediction.  [31] introduces the concept of dropout training as an approximate Bayesian inference method for deep Gaussian processes, providing a robust tool for modeling model uncertainty.  [32] distinguishes between two primary types of uncertainty: data uncertainty (aleatoric uncertainty) and model uncertainty (epistemic uncertainty). They have proposed a Bayesian deep learning framework that integrates these uncertainties, which has been applied to tasks such as semantic segmentation and deep regression. HiVT [33] is proposed to enhance the accuracy and efficiency of motion prediction by modeling interactions between multiple agents through a hierarchical structure.  [34] extends various online map estimation methods by incorporating uncertainty estimation and strengthening the role of map information in trajectory prediction.
III System Overview and Formulation
III-A System Overview of ReNiL
@@
The workflow of ReNiL is shown in Fig. 1. The proposed ReNiL system consists of three main stages: orientation alignment, any-scale Laplace estimator, and Bayesian process. 1) The orientation alignment stage continuously collects accelerations, angular rates, and magnetic fields in the background, dynamically aligning the smartphone coordinate system with the navigation coordinate system (e.g. ENU). It updates the orientation of device through a motion-aware orientation filter. 2) The any-scale Laplace estimator handles aligned IMU sequences of any scale by extracting local patch features and further building contextual features. These features are combined via scale pooling into unified global features, which are subsequently used to estimate both the displacement and the uncertainty parameter modeled by a Laplace distribution. 3) The Bayesian process stage incorporates these displacement and uncertainty estimations into a Bayesian inference framework, sequentially refining the positioning result at each IPDP.
III-B Formulation for Bayesian IPDP Localization Task
@@
Step I: This step is the orientation alignment stage. It ensures the directional consistency of the network’s relative displacement output. A pedestrian inertial localization task generally involves two right-handed Cartesian coordinate systems: the navigation coordinate system and the device coordinate system. To avoid numerical issues and singularities, we use a quaternion, denoted by $q=w+xi+yj+zk$ to describe the rotation from the device coordinate system to the navigation coordinate system. For a timestamp measurement $(a,b,c)$ in the device coordinate system, and its pure quaternion $X=ai+bj+ck$, quaternion multiplication $\otimes$ is used to rotate the vector to the navigation coordinate system, as expressed by $X^{g}=q^{-1}\otimes X\otimes q$. For the inertial task, this process involves rotating the nine-axis measurements consisting of the accelerometer ($A\in\mathcal{R}^{3}$), gyroscope ($G\in\mathcal{R}^{3}$), and magnetometer ($M\in\mathcal{R}^{3}$) to the navigation coordinate system:
@@
And then update the quaternion $q$ to a new quaternion $q_{new}$ by orientation filter $f_{o}(\cdot;\cdot)$:
$q_{new}=f_{o}(\begin{pmatrix}A\ G\ M\end{pmatrix},q)$ (2)
Step II: This step constitutes the core of the Bayesian IPDPs pedestrian neural network task. We consider a sequence of aligned inertial measurements $x_{i}^{g}=[A^{g}_{i},G^{g}_{i}]$ from the IPDP $t_{n-1}$ to $t_{n}$:
$X^{g}_{t_{n}}=\left\{x^{g}_{t_{n-1}},\ldots,x^{g}_{t_{n}}\right\},\quad{x}^{g}_{t}\in\mathcal{R}^{6}$ (3)
The goal of the localization task at IPDPs is to estimate the relative displacement at specific demand points instead of continuously estimating the trajectory. To quantify uncertainty at each demand point, the ground truth displacement $\Delta p_{t_{n}}\in\mathcal{R}^{2}$ can be modeled probabilistically via a Laplace distribution:
$\Delta p_{t_{n}}\sim\operatorname{Laplace}\left(\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\right),\quad\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\in\mathcal{R}^{2}$ (4)
where $\Delta\hat{p}_{t_{n}}$ is the predicted displacement, $\hat{b}_{t_{n}}$ is the estimated uncertainty scale parameter at the IPDPs.
@@
$(\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}})=f\left(X^{g}_{t_{n}};\theta\right)$ (5)
@@
Step III: This step describes the localization task of multiple consecutive demand points as a Bayesian localization inference framework. Assume that there is a series of IPDPs, expressed as:
${T}_{d}=\left\{t_{0},t_{1},\ldots,t_{n}\right\},\quad 0\leq t_{0}<t_{1}<\cdots<t_{n}\leq T$ (6)
@@
$\mathbf{p}\left({p}_{t_{n}}\right)=\int\mathbf{p}\left({p}_{t_{n}}\mid{p}_{t_{n-1}},\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\right)\mathbf{p}\left({p}_{t_{n-1}}\right)d{p}_{t_{n-1}}$ (7)
@@
IV Relative Neural Inertial Locator
IV-A Orientation Alignment

### v3

# ReNiL: Event-Driven Pedestrian Bayesian Localization Using IMU for Real-World Applications

@@ Abstract
Pedestrian inertial localization plays a crucial role in mobile and Internet of Things (IoT) services, offering ubiquitous positioning without requiring external infrastructure. However, existing deep-learning-based pedestrian inertial localization methods rely on fixed sliding-window integration. These methods lack flexibility in adapting to varying motion scales and cadences, making them difficult to transfer across different devices, users, and motion patterns. Additionally, they cannot estimate uncertainties and errors consistent with the data source. This limits their utility in real-world mobile scenarios. In this paper, we propose Relative Neural Inertial Locator (ReNiL), a novel deep learning-based event-driven Bayesian inference framework for accurate, efficient, and uncertainty-aware pedestrian localization. ReNiL introduces the concept of Inertial Positioning Demand Points (IPDPs) and focuses estimation on contextually meaningful points rather than continuous dense tracking. It supports inference on inertial measurement unit (IMU) sequences at any scale, enabling flexible positioning cadences aligned with application needs, thus saving significant computational overhead. ReNiL integrates a motion-aware orientation filter with an Any-Scale Laplace Estimator (ASLE), a dual-task deep network that combines patch-based self-supervision and Bayesian regression. By parameterizing displacement with a Laplace distribution, ReNiL yields a high-precision estimate of the displacement as well as its Euclidean-homogeneous uncertainty estimate. Then a Bayesian inference chain connects successive IPDPs and other positioning sources to form an efficient and unified localization process. Extensive experiments on RoNIN-ds and our new WUDataset, which feature complex indoor/outdoor motion from 28 participants using multiple different types of devices, have demonstrated that ReNiL achieves state-of-the-art position estimation accuracy and uncertainty consistency, outperforms TLIO, CTIN, iMoT, and RoNIN variants. Application studies have confirmed ReNiL’s ubiquity, robustness and usability for mobile and IoT localization tasks, establishing it as a scalable, uncertainty-aware foundation for next-generation mobile positioning systems.
@@
Keywords: Pedestrian localization, IMU, event-driven, Bayesian process, deep learning
1 Introduction
The ubiquitous use of smart mobile devices has increased the demand for location-based services. Pedestrian localization is an important technique that supports indoor navigation, AR/VR interaction, and safety-critical guidance. The goal of a pedestrian localization system is to deliver accurate position estimates together with well calibrated error characterization at the time decisions are required. Existing approaches fall into two classes: (i) infrastructure- or map-dependent methods, including radio fingerprinting and ranging with WiFi, BLE, UWB, or RFID [shang2022overview, kriz2016improving, gezici2009position, bahl2000radar, xu2023principle, 10.1145/3699792, 10.1145/3678554, 10.1145/3214287] and vision- or LiDAR-based odometry (VIO/LIO), which typically requires prebuilt maps or rich visual texture [scaramuzza2011visual, zhang2014loam, 10.1145/3161409, 10.1145/3757940.3757942, 10.1145/3669721.3669749]; (ii) inertial methods that operate without external signals. Inertial methods use an inertial navigation system (INS) to integrate a specific force and angular rate measured by an inertial measurement unit (IMU).
This process has gain traction in recent years as the IMU propagates the state over time without infrastructure or prior maps, improving privacy, and simplifies deployment. In addition, the adoption of deep learning for consumer-grade MEMS IMUs has further improved IMU-only localization. Compared to classical pedestrian dead-reckoning (PDR) and strapdown INS with ZUPT or magnetometer updates [titterton2004strapdown, anguita2013public], data-driven methods show that convolutional-, recurrent-, and attention-based models infer relative displacement from IMU sequences and surpass traditional smartphone baselines [herath2020ronin, liu2020tlio, wald2019rio, 8937008, sun2021idol, rao2022ctin, nguyen2025imot, 10.1145/3534594]. Probabilistic models further estimate heteroscedastic uncertainty. These models are typically under a Gaussian assumption and trained by minimizing the negative log-likelihood, which facilitates downstream fusion [liu2020tlio, sun2021idol, 8937008, rao2022ctin]. Moreover, IMU can be combined with external sources, such as GNSS, radio, vision, or LiDAR using filtering or factor-graph formulations, for higher accuracy  [al2021map, magsi2024comparison, lee2021indoor, thio2021fusing, 10.1145/3722570.3726876, 10.1145/3662009.3662018, 10.1145/3214270, 10.1145/3328918, 10.1145/3569467].
While great progress has been made facilitating higher performance of pedestrian localization, most deep-learning based inertial methods use a fixed sliding window and directly regress an IMU segment to the displacement or velocity of uniformly spaced trajectory points. This design generates acceptable performance in offline evaluations, but faces three challenges in real-world, online deployment:
  • Cadence mismatch and compute waste. Real-world applications are event driven (e.g., doorway crossings, turns, stairs transitions, arrivals at points of interest). These events are not triggered in fixed time windows. Fixed-rate inferences are excessive and creates unnecessarily dense samples that accumulate drift. The inference cadence does not match decision times, and increases device resource usage.
  • Uncertainty miscalibration for fusion. Many networks either omit uncertainty or assume a single Gaussian model. Nonstationary motion and sensor jitter, where heavy-tailed errors are common. Empirical coverage and negative log-likelihood diverge from the residual statistics. As a result, outputs lack a reliable scale in a shared Euclidean state space, which hinders principled fusion with other sources and degrades subsequent filtering and smoothing.
  • Single-scale bias and poor transfer. A fixed window cannot capture both short, high-dynamic segments and long, slowly varying context. The resulting scale mismatch reduces transfer across users, carry patterns, and devices, and yields brittle performance in heterogeneous environments.
@@ 1 Introduction
To address the challenges above, we propose ReNiL, a deep learning-based Bayesian inference framework for inertial localization. To eliminate excessive estimation of relative displacement and uncertainty, we define an innovative concept called inertial positioning demand points (IPDPs). ReNiL only compute the estimates at IPDPs. ReNiL also incorporates a motion-aware orientation filter specifically designed for pedestrian dynamics, enhancing the alignment between IMU measurements and the navigation coordinate system. To support IMU sequences of any scale and enable uncertainty estimation, it employs a dual-task deep network, Any-Scale Laplace Estimator (ASLE), which combines patch-wise self-supervision with Bayesian regression. Furthermore, ReNiL establishes a principled Bayesian inference pipeline that forms a continuous chain of displacement and uncertainty estimates across IPDPs.
@@
When the user starts using the inertial service, the system collects IMU data to adaptively estimate the device orientation and gradually align the navigation coordinate system with the ENU coordinate system. ReNiL first aligns the entire IMU sequence with the navigation coordinate system based on the device orientation at each timestamp and then patches the data according to a given patch length. The patched data are processed in parallel by an input module and a feature extractor to extract motion features. To improve temporal correlation, the features are reconstructed into robust contextual motion features by a contextual builder. A scale-pooling multi-head self-attention mechanism is designed to build unified global features and contextual features, allowing it to regress with different scales. The self-supervision patch method is used to further improve the noise resistance of the model and the contextual learning capabilities. To handle potential aleatoric uncertainty in the data and improve the numerical stability of regression, displacement and its uncertainty estimates are parameterized using the Laplace distribution [laplace1774memoire]. Additionally, ReNiL provides a general process based on Bayesian inference for pure IPDP tasks or tasks loosely coupled with external positioning sources.
@@
Extensive experiments on the open-source RoNIN-ds and our own WUDataset have demonstrated that ReNiL achieves state-of-the-art position estimation accuracy and uncertainty consistency, surpassing all advanced models. We have verified the validity of parameterizing the displacements with the Laplace distribution and validated its Euclidean homogeneity. We have analyzed the computational overhead of ReNiL and several other models, and verified that ReNiL can greatly reduce the parameters and overhead of floating-point operations. Finally, we have conducted a series of practical application experiments, demonstrating the high availability of ReNiL in real-world scenarios by combining the Bayesian process.
@@
  • IPDPs task formulation. To the best of our knowledge, this is the first paper that defines pedestrian localization as an on-demand task with IPDPs, instead of using fixed-rate estimation.
@@
  • Comprehensive validation. Experiments on RoNIN-ds and our WUDataset show that ReNiL achieves state-of-the-art position estimation accuracy, Euclidean homogeneous uncertainty, and markedly reduces FLOPs.
2 Related Work
2.1 Pedestrian Dead Reckoning
Pedestrian Dead-Reckoning (PDR) estimates a pedestrian’s position by integrating step detection, stride length estimation, and heading direction from inertial sensors [brajdic2013walk]. Traditional PDR methods are highly dependent on heuristic models and sensor calibration, which may suffer from cumulative drift and noise. Recent work has focused on improving robustness by using map information [kang2021indoor] or combining additional signals such as magnetometers and barometers [liu2024reliable]. However, conventional PDR still faces challenges in complex indoor environments with magnetic disturbances and sensor biases [hou2020pedestrian].
2.2 Deep Learning for Inertial Navigation
Deep learning has been increasingly applied to inertial navigation to address the limitations of traditional algorithms. Data-driven methods learn representations from raw or aligned IMU data to directly estimate displacement. RNNs, especially LSTM networks, have shown effectiveness in modeling temporal dependencies in IMU sequences [8937008, sun2021idol]. CNN-based architectures have been explored to effectively suppress short-term drift in sensor data [herath2020ronin, liu2020tlio, wald2019rio]. Attention-based architectures have also been explored to capture long-range correlations in sensor data [rao2022ctin, nguyen2025imot]. These methods reduce the need for manual calibration and demonstrate improved accuracy.
2.3 Uncertainty Estimation
In the context of deep learning, uncertainty modeling has emerged as a critical technique to improve model reliability and robustness, particularly in domains such as computer vision and trajectory prediction. [gal2016dropout] introduces the concept of dropout training as an approximate Bayesian inference method for deep Gaussian processes, providing a robust tool for modeling model uncertainty. [kendall2017uncertainties] distinguishes between two primary types of uncertainty: data uncertainty (aleatoric uncertainty) and model uncertainty (epistemic uncertainty). They have proposed a Bayesian deep learning framework that integrates these uncertainties, which has been applied to tasks such as semantic segmentation and deep regression. HiVT [zhou2022hivt] is proposed to enhance the accuracy and efficiency of motion prediction by modeling interactions between multiple agents through a hierarchical structure. [gu2024producing] extends various online map estimation methods by incorporating uncertainty estimation and strengthening the role of map information in trajectory prediction.
3 System Overview and Formulation
@@
3.1 System Overview of ReNiL
@@
The workflow of ReNiL is shown in Fig. 1. The proposed ReNiL system consists of three main stages: orientation alignment, any-scale Laplace estimator, and Bayesian process. (1) The orientation alignment stage continuously collects accelerations, angular rates, and magnetic fields in the background, dynamically aligning the smartphone coordinate system with the navigation coordinate system (e.g., ENU). It updates the orientation of the device through a motion-aware orientation filter. (2) The any-scale Laplace estimator handles aligned IMU sequences of any scale by extracting local patch features and further building contextual features. These features are combined via scale pooling into unified global features, which are subsequently used to estimate both the displacement and the uncertainty parameter modeled by a Laplace distribution. (3) The Bayesian process stage incorporates these displacement and uncertainty estimations into a Bayesian inference framework, sequentially refining the positioning result at each IPDP.
3.2 Formulation for Bayesian IPDP Localization Task
@@
Step I: This step is the orientation alignment stage. It ensures the directional consistency of the network’s relative displacement output. A pedestrian inertial localization task generally involves two right-handed Cartesian coordinate systems: the navigation coordinate system and the device coordinate system. To avoid numerical issues and singularities, we use a quaternion, denoted by $q=w+xi+yj+zk$, to describe the rotation from the device coordinate system to the navigation coordinate system. For a timestamp measurement $(a,b,c)$ in the device coordinate system, and its pure quaternion $X=ai+bj+ck$, quaternion multiplication $\otimes$ is used to rotate the vector to the navigation coordinate system, as expressed by
@@
Then the quaternion $q$ is updated to a new quaternion $q_{\text{new}}$ by the orientation filter $f_{o}(\cdot;\cdot)$:
$q_{\text{new}}=f_{o}\!\left(\begin{pmatrix}A\ G\ M\end{pmatrix},q\right).$ (2)
Step II: This step constitutes the core of the Bayesian IPDPs pedestrian neural network task. We consider a sequence of aligned inertial measurements $x_{i}^{g}=[A_{i}^{g},G_{i}^{g}]$ from the IPDP $t_{n-1}$ to $t_{n}$:
$X^{g}_{t_{n}}=\left\{x^{g}_{t_{n-1}},\ldots,x^{g}_{t_{n}}\right\},\quad x^{g}_{t}\in\mathcal{R}^{6}.$ (3)
The goal of the localization task at IPDPs is to estimate the relative displacement at specific demand points instead of continuously estimating the trajectory. To quantify uncertainty at each demand point, the ground-truth displacement $\Delta p_{t_{n}}\in\mathcal{R}^{2}$ can be modeled probabilistically via a Laplace distribution:
$\Delta p_{t_{n}}\sim\operatorname{Laplace}\!\left(\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\right),\quad\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\in\mathcal{R}^{2},$ (4)
where $\Delta\hat{p}_{t_{n}}$ is the predicted displacement, and $\hat{b}_{t_{n}}$ is the estimated uncertainty scale parameter at the IPDPs.
@@
$(\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}})=f\!\left(X^{g}_{t_{n}};\theta\right),$ (5)
@@
Step III: This step describes the localization task of multiple consecutive demand points as a Bayesian localization inference framework. Assume that there is a series of IPDPs, expressed as
$T_{d}=\left\{t_{0},t_{1},\ldots,t_{n}\right\},\quad 0\leq t_{0}<t_{1}<\cdots<t_{n}\leq T.$ (6)
@@
$\mathbf{p}\!\left(p_{t_{n}}\right)=\int\mathbf{p}\!\left(p_{t_{n}}\mid p_{t_{n-1}},\Delta\hat{p}_{t_{n}},\hat{b}_{t_{n}}\right)\mathbf{p}\!\left(p_{t_{n-1}}\right)dp_{t_{n-1}}.$ (7)
@@
4 Relative Neural Inertial Locator
4.1 Orientation Alignment